OpenStudy (anonymous):

PLEASE HELP, easy. What are the exact solutions of x^2-5x-7=0 ?

OpenStudy (anonymous):

This does not factor.

OpenStudy (anonymous):

does no factor

OpenStudy (anonymous):

I know, u have to use the quadratic formula

OpenStudy (jiteshmeghwal9):

use formula\[-b \pm \sqrt{b^2-4ac} \over 2a\]

OpenStudy (anonymous):

But I mess up on it I can't do it!!

OpenStudy (anonymous):

Cool, use the formula then =)

OpenStudy (anonymous):

Alright I'll try :)

OpenStudy (anonymous):

you need to know a,b, and c a=1 b=5 c=7 Now substitute this into the quadratic formula and solve

OpenStudy (anonymous):

Okay thanks guys

OpenStudy (anonymous):

@aussy123 b=-5 and c=-7

OpenStudy (anonymous):

Oh they are negative?

OpenStudy (anonymous):

Yes. b and c are negative.

OpenStudy (anonymous):

Thnx

OpenStudy (jiteshmeghwal9):

so,\[5 \pm \sqrt{25+28} \over 2\]after solving these make two roots on as\[\alpha=\frac{-b + \sqrt{b^2-4ac}}{2a}\] & second as\[\beta=\frac{-b -\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

@SmartWish Thanks for the correction

OpenStudy (anonymous):

So wait

OpenStudy (anonymous):

You get 53 over 2

OpenStudy (anonymous):

@aussy123

OpenStudy (anonymous):

Im checking now

OpenStudy (anonymous):

use quasratic equation

OpenStudy (anonymous):

Oh got it

OpenStudy (anonymous):

use quadratic equation -b ± √b²-4ac x=------------------ 2a a=1 b=-5 c=-7 put them in above equation

OpenStudy (anonymous):

\[\frac{5\pm \sqrt{53}}{2}\]

OpenStudy (anonymous):

there was a typo mistake correct is as below

OpenStudy (anonymous):

Thanks guys appreciate it